# muography; non-destructive testing; machine learning; image processing; upsampling; semantic segmentation; simulation; cosmic rays

### v1

@@ 4.3 Future Directions: Model Refinement and Defect Detection Applications
The dataset used to train both models covered a wide range of sampling rates. Upsampling performance on the lowest generated image data was much better than anticipated, whereas performance at sampling rates greater than 50–60 days yielded diminishing returns as well as indications from convergence between input and outputs metric scores. Reducing the maximum sampling rate would drastically decrease required computation time for simulations and PoCA algorithm processing. The high performance of the upsampling model on 1-day sampling rates, especially shown by the visual examples, indicates that the limits of the model should be further pushed to provide results for sampling rates on the order of hours.
Figure 5: A single XZ-plane vertical image slice for different sampling rates: (a) one day, (b) five days, (c) 10 days, (d) 20 days, (e) 40 days, (f) 60 days, (g) 80 days, (h) ground truth (100-day image for the top two panels, geometry truth for the bottom two panels). These eight image versions are displayed as raw input (first), upsampled (second down), segmented (third down), and upsampled and segmented (bottom). Lilac, blue and red indicate concrete, rebar and tendon ducts respectively.
@@
Despite using segmentation purely as an analysis tool, its ability to resolve true object boundaries and therefore decrease smearing effects is a significant result. Furthermore, it was noted that segmentation performance on the upsampled images was limited, despite there being clear objects present. This may be a symptom of using a two-model approach, as the outputs of the upsampling model $\hat{y}$ are similar but not identical to the input $y$ used to train the segmentation model. In order to achieve the most accurate outputs, without the complexity of two models, we can simplify the task to a one-model approach which uses the true object geometries to generate a ground truth — whether that be a segmentation map of labels, or of material-specific values such as density or radiation length.
@@
The limited receptive field of convolutional kernels restricted the model to having a localised receptive field — prevalent in the non-uniformity of rebar grids or straight edges of the segmented images. While this issue can be circumvented through the use of hierarchical dilated convolutions Chen et al. (2018), allowing for a slight increase in context size, the most popular solution is to use attention mechanisms Vaswani et al. (2017). Attention is a more recent and increasingly prominent methodology for model construction, powering state-of-the-art models in both language and vision tasks. It uses a tokenisation system that splits images into patches, which are flattened and processed as individual tokens. These tokens are positionally embedded to preserve spatial relationships, enabling the model to account for location dependence. Unlike convolutions, which have a limited receptive field that grows with depth, attention mechanisms inherently consider relationships between all tokens, providing a global context at every layer. This ability to capture long-range dependencies and spatial relationships makes attention a powerful alternative for vision tasks.
@@
This study has looked at the upsampling and identification of objects commonly found within concrete interiors. However, the identification of objects solves only part of the problem pertaining the characterisation of built infrastructure. The next step is to alter the model’s objective to look towards not only identifying object location and material, but to also accurately identify defects — such as tendon duct strand corrosion, honeycombing and air voids.
@@
Despite the general agreement of Geant4 simulations and real data for muography, it is important to verify these models on real-world data. Therefore, testing of the models on data gathered from Niederleithinger et al. Niederleithinger et al. (2021), as well as planned future muography scans by Lynkeos Technology Ltd, will be important for validation of the results of this study. Additionally, the use of real data integrated into the simulated datasets will aid the model in adapting to real-world conditions.
@@ 5 Conclusions
In this study, we have demonstrated the successful application of deep learning techniques to address two significant challenges in muon scattering tomography: reducing required sampling time and improving feature identification in concrete structures. Our cWGAN-GP model achieved remarkable upsampling performance, with 1-day sampled images being enhanced to match the quality metrics of 21 baseline images for SSIM and 31 baseline images for PSNR. This represents a significant reduction in required sampling time for achieving high-quality muography images. To evaluate the model’s performance beyond traditional metrics, we developed an approach using semantic segmentation as an assessment tool. This method revealed that the upsampling model’s effectiveness varies significantly across different structural features, with particularly strong performance in enhancing the visibility of tendon ducts and moderate improvements in rebar grid detection. The segmentation model also demonstrated an unexpected capability to differentiate between real features and shadowing artefacts caused by a low vertical resolution, thus addressing the inverse imaging problem inherent to muography. Whilst our results are promising, key challenges remain. The model’s performance on air void detection requires improvement, likely through the implementation of class-weighted loss functions to address class imbalance. Additionally, the current approach of processing 2D slices independently limits the model’s ability to utilise full 3D contextual information, suggesting potential improvements through multi-plane or patch-based processing methods. The transition from simulation to real-world application represents the next critical step in this research. Future work will focus on validating these models with experimental data and extending the segmentation capabilities to identify specific structural defects, such as tendon duct corrosion and concrete honeycombing. These advancements could significantly impact the practical application of muography in infrastructure inspection, potentially reducing inspection times while improving the accuracy of defect detection.
@@
Conceptualization, W.O., D.M., G.Y. and S.G.; methodology, W.O.; software, W.O. and G.Y.; validation, W.O.; formal analysis, W.O.; investigation, W.O.; resources, W.O.; data curation, W.O.; writing —- original draft preparation, W.O.; writing—-review and editing, D.M. and G.Y.; visualisation, W.O.; supervision, D.M., G.Y. and S.G.; project administration, D.M.; funding acquisition, D.M. All authors have read and agreed to the published version of the manuscript.
@@
This research was funded under a Ph.D. studentship by Lynkeos technology Ltd. and the College of Science and Engineering at the University of Glasgow.
@@
Data available upon request.

### v2

@@ 4.3 Future Directions: Model Refinement and Defect Detection Applications
The dataset used to train both models covered a wide range of sampling times. The upsampling performance on the lowest generated image data was much better than anticipated, whereas performance at equivalent sampling times greater than 50–60 days yielded diminishing returns as well as indications of convergence between the input and output metric scores. Reducing the maximum sampling time would drastically decrease the required computation time for simulations and PoCA algorithm processing. The high performance of the upsampling model on a 1-day equivalent sampling time, especially shown by the visual examples, indicates that the limits of the model should be further pushed to provide results for equivalent sampling times on the order of hours.
@@
Despite using segmentation purely as an analysis tool, its ability to resolve true object boundaries and therefore decrease smearing effects is a significant result. Furthermore, it was noted that the segmentation performance on the upsampled images was limited, despite there being clear objects present. This may be a symptom of using a two-model approach, as the outputs of the upsampling model $\hat{y}$ are similar but not identical to the input $y$ used to train the segmentation model. In order to achieve the most accurate outputs, without the complexity of two models, we can simplify the task to a one-model approach which uses the true object geometries to generate a ground truth—whether that be a segmentation map of labels or of material-specific values such as density or radiation length.
@@
The limited receptive field of convolutional kernels restricted the model to having a localised receptive field—prevalent in the non-uniformity of the rebar grids or straight edges of the segmented images. While this issue can be circumvented through the use of hierarchical dilated convolutions Chen et al. (2018), allowing for a slight increase in context size, the most popular solution is to use attention mechanisms Vaswani et al. (2017). Attention is a more recent and increasingly prominent methodology for model construction, powering state-of-the-art models in both language and vision tasks. It uses a tokenization system that splits images into patches, which are flattened and processed as individual tokens. These tokens are positionally embedded to preserve spatial relationships, enabling the model to account for location dependence. Unlike convolutions, which have a limited receptive field that grows with depth, attention mechanisms inherently consider relationships between all tokens, providing a global context at every layer. This ability to capture long-range dependencies and spatial relationships makes attention a powerful alternative for vision tasks.
While the concrete samples used to train the models are of realistic design, the dataset poorly reflects a wide variety real-world scenarios which include more varied object orientations, concrete thicknesses, and detector orientations. Therefore, future dataset generation should aim to vary these parameters to allow the model to generalise to different scenarios.
@@
This study has looked at the upsampling and identification of objects commonly found within concrete interiors. However, the identification of objects solves only part of the problem pertaining to the characterization of built infrastructure. The next step is to alter the model’s objective to look toward not only identifying object location and material but to also accurately identify defects—such as tendon duct strand corrosion, honeycombing, and air voids.
@@
Despite the general agreement of Geant4 simulations and real data for muography, it is important to verify these models on real-world data. Therefore, testing of the models on data gathered from Niederleithinger et al. Niederleithinger et al. (2021), as well as planned future muography scans by Lynkeos Technology Ltd, will be important for validation of the results of this study.
Figure 5: A single X–Z plane vertical image slice for different equivalent sampling times: (a) one day, (b) five days, (c) 10 days, (d) 20 days, (e) 40 days, (f) 60 days, (g) 80 days, (h) ground truth (100-day image for the top two panels, geometry truth for the bottom two panels). These eight image versions
are displayed as raw input (first), upsampled (second down), segmented (third down), and upsampled and segmented (bottom). Lilac, blue, and red indicate concrete, rebar, and tendon ducts, respectively.
@@ 5 Conclusions
In this study, we have demonstrated the successful application of deep learning techniques to address two significant challenges in muon scattering tomography: reducing required sampling times and improving feature identification in concrete structures. Our cWGAN-GP model achieved remarkable upsampling performance, with 1-day sampled images being enhanced to match the quality metrics of 21 baseline images for the SSIM and 31 baseline images for the PSNR. This represents a significant reduction in the required sampling times for achieving high-quality muography images. To evaluate the model’s performance beyond traditional metrics, we developed an approach using semantic segmentation as an assessment tool. This method revealed that the upsampling model’s effectiveness varies significantly across different structural features, with particularly strong performance in enhancing the visibility of tendon ducts and moderate improvements in rebar grid detection. The segmentation model also demonstrated an unexpected capability to differentiate between real features and shadowing artefacts caused by a low vertical resolution, thus addressing the inverse imaging problem inherent to muography. While our results are promising, key challenges remain. The model’s performance on air void detection requires improvement, likely through the implementation of class-weighted loss functions to address class imbalance. Additionally, the current approach of processing 2D slices independently limits the model’s ability to utilise full 3D contextual information, suggesting potential improvements through multi-plane or patch-based processing methods. The transition from simulation to real-world application represents the next critical step in this research. Future work will focus on generalizing the models to a range of realistic imaging scenarios, as well as validating these models with experimental data and extending the segmentation capabilities to identify specific structural defects, such as tendon duct corrosion and concrete honeycombing. These advancements could significantly impact the practical application of muography in infrastructure inspection, potentially reducing inspection times while improving the accuracy of defect detection.
@@
Conceptualization, W.O., D.M., G.Y., and S.G.; methodology, W.O.; software, W.O. and G.Y.; validation, W.O.; formal analysis, W.O.; investigation, W.O.; resources, W.O.; data curation, W.O.; writing—-original draft preparation, W.O.; writing—-review and editing, D.M. and G.Y.; visualization, W.O.; supervision, D.M., G.Y., and S.G.; project administration, D.M.; funding acquisition, D.M. All authors have read and agreed to the published version of the manuscript.
@@
This research was funded through an agreement between Lynkeos Technology Ltd. and the College of Science and Engineering at the University of Glasgow.
@@
The data are available upon request.